OpenStudy (anonymous):

5x^3y^2+13x^4y^3-17x^3y^4 degree

OpenStudy (mathstudent55):

You need the degree of the polynomial?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i know its not 7

OpenStudy (mathstudent55):

A monomial is a single number or a number multiplied by any number of variables. A sum of monomials is a polynomial. The degree of a monomial is the sum of all the exponents of the variables of that monomial. The degree of a ploynomial is the same as the degree of the monomial with highest degree

OpenStudy (mathstudent55):

Why do you say you know it's not 7?

OpenStudy (anonymous):

Because I thought you would add the polynomials of each set of numbers (ex. 5x^3y^2), but that isn't correct the highest would be 7 for 3+4, but that is incorrect

OpenStudy (mathstudent55):

Add the exponents of each term: 3 + 2 = 5 4 + 3 = 7 3 + 4 = 7 Highest degree in a term = 7 (both the second and third terms) Degree of polynomial = 7

OpenStudy (mathstudent55):

7 is correct for the degree of the polynomial. The degree of each term is: 5, 7, 7, respectively.

OpenStudy (anonymous):

thats what i thought, but the program i entered it into said 7 was incorrect

OpenStudy (mathstudent55):

Are you sure you copied the problem correctly?

OpenStudy (anonymous):

Yes, I already checked. It might just be a system error.

OpenStudy (anonymous):

Thanks for your help.

OpenStudy (mathstudent55):

You're welcome.